Write an equation in slope intercept form for the line that satisfies the following condition. slope (7/13) and passes through (6,-15)
Answer by mananth(16949)

Write an equation in slope intercept form for the line that satisfies the following condition. slope (7/13) and passes through (6,-15)
...
y= mx +b
plug values of x.y & slope
-15=7/13*6+b
-15=42/13+b
b=-15-(42/13)
b=-237/13
..
y=(7/13)x-(237/13)